# Revisiting the Reliability of Language Models in Instruction-Following

### v1

@@ Abstract
Jianshuo Dong1, Yutong Zhang1, Yan Liu2, Zhenyu Zhong2, Tao Wei2, Chao Zhang1, Han Qiu1∗ 1Tsinghua University, China. 2Ant Group, China.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable abilities to follow natural language instructions (ouyang2022instructgpt). The instruction-following ability is foundational for enabling faithful user interaction (zhou2023instruction-ifeval-verification-instruction-following-evaluation-25-types; lior2025wildifeval; zhang2024cfbench-comprehensive-constraints-following-benchmark-for-llms), reliable agent behavior (zhang2025which-multi-agent-system-failure-attribution), and reduced harmful outcomes (ruan2024identifying-tool-emu-simulate-tools-to-test-agent-safety). Evaluating this ability is essential for building reliable, trustworthy AI systems. Consequently, numerous benchmarks have been proposed to measure instruction-following performance across diverse task types and constraint categories (see Appendix˜A for a survey of 36 benchmarks).
@@
As illustrated in Figure˜1, the instruction-following ability of LLMs has improved rapidly across generations. Recent LLMs achieve near-ceiling performance on widely adopted benchmarks like IFEval (zhou2023instruction-ifeval-verification-instruction-following-evaluation-25-types), with GPT-5 reaching an accuracy as high as 95.9%. However, growing evidence reveals that model performance can be highly sensitive to prompt wording (sclar2024quantifying-how-i-learned-to-start-worrying-about-prompt-formatting-classification-tasks; mizrahi2024state-multi-prompt-llm-evaluation; cao2024worst-prompt-performance-of-llms), and that overfitting to benchmark test cases is possible (zhang2024a-benchmark-over-fitting). Thus, questions remain about how well benchmark accuracy translates to real-world reliability.
@@
To operationalize this, we propose a new metric, reliable@k, which quantifies whether an LLM can simultaneously solve a set of cousin prompts. We further design an automated pipeline for generating high-quality cousin prompts using three data augmentation strategies: rephrasing, distractor addition, and task/constraint reconfiguration. This is coupled with a code-assisted validity checker. Building on IFEval (zhou2023instruction-ifeval-verification-instruction-following-evaluation-25-types), we leverage the pipeline to construct an extended benchmark, IFEval++ ***We name it to show respect for the renowned IFEval benchmark, as we position IFEval++ as a successor of IFEval, similar to how C++ relates to C. , which consists of 541 test cases, each comprising 10 cousin prompts. This provides a testbed for assessing nuance-oriented reliability of LLMs in a systematic way.
@@ 2.2 Scope of This Work
Data Source & Constraint Types. We choose IFEval (zhou2023instruction-ifeval-verification-instruction-following-evaluation-25-types) as our main testbed, since SOTA LLMs have already achieved strong performance on it (e.g., OpenAI’s GPT-5 reaches an accuracy of 95.9% as measured in Table˜1), making it a meaningful setting for discussing reliability and the challenging reliable@k metric. This benchmark focuses on single-turn tasks with 25 types of code-verifiable format constraints, such as “Answer using at most {num_words} words.” Reliably tackling such simple yet foundational instructions is a necessary condition for addressing more complex, multi-turn, and multi-modal instruction-following settings. See Section˜B.1 for the 25 constraints.
@@ 3.1 Data Augmentation Methods
Type 2. Distractor Addition. The augmenter appends each prompt with a distractor constraint that does not interfere with the satisfaction of the original constraints (see Section˜E.2). The augmenter is instructed to ensure that each distractor constraint remains compatible with the original prompt. In this setting, responses continue to be evaluated against the original constraints, with the added constraint functioning purely as a distractor, in contrast to prior work on multi-constraint composition (jiang2024followbench-multi-constraints-composition-and-considering-decoding-methods; wen2024benchmarking-complexbench-multiple-constraints-composition). This setup examines whether LLMs can integrate additional constraints in a plug-and-play manner.
@@ 3.2 Code-Assisted Validity Check
The checker is designed with an emphasis on high recall: its primary objective is to flag any potentially erroneous cases, even at the cost of over-detection. This strategy allows us to curate a high-quality evaluation set by filtering candidates against the checker. Inspired by chen2023program-of-thoughts-prompting, our central design is to embed the implementations of evaluation functions directly into the checker’s prompt, clearly specifying how the evaluation operates (see Section˜E.2). Empirically, this approach enables more deterministic and robust verification. The checker is implemented in a unified way to monitor all three types of augmented test cases.
@@
To further assess robustness, we evaluated it on 3×1,000=3,0003$\times$ 1,000=3,000 additional flawed cases, maintaining a recall of 99.9% (2,997/3,000); the few misses were due to unavoidable hallucinations (sun2025detection-hallucination-of-lrms). The checker also achieved 90.0% precision on 541 unrevised promptsMany flagged cases were indeed flawed or ambiguous. Necessary corrections are detailed in Section B.3, and the revised dataset is used for subsequent experiments. . Overall, the checker rigorously detects invalid prompts with minimal valid loss. Moreover, since the augmenter is also explicitly guided to avoid introducing flaws, the two-layer mechanism minimizes the risk of flaws arising from the augmentation process.
@@ 4.1 Experimental Setup
Evaluation Methods. Following the practices of DeepSeek-V3 (liu2024deepseek-v3), Qwen3 (qwen2025qwen3technicalreport), and Kimi K2 (team2025kimi-k2), we use the prompt strict evaluation mode of IFEval. This requires LLMs to satisfy all requirements specified in the prompts. Based on this success signal, we compute the reliable@k metric.
@@ 4.2 Main Results with IFEval++
$\bullet$ Augmentation Type. Models exhibit varying robustness to the three types of cousin prompts. Most models can effectively handle rephrased cousin prompts, which share the greatest similarity with the original ones. This result is expected, as rephrased prompts differ only in wording (understanding the instructions) while requesting the same underlying response (executing the instructions). In contrast, cousin prompts that include additional distractor constraints or C/T reconfiguration introduce greater complexity. Their corresponding reliable@4 scores are typically lower than those of rephrased prompts. We hypothesize that, although distractor constraints are designed to remain compatible with the original instructions, and C/T reconfiguration introduces only slight variations in the underlying requirements, both inevitably increase the difficulty of response planning and execution (dong2025emergent-planning-pre-generation-probing).
@@ 5.1 Predicting the Selective Following
If we can predict how the LLM reacts to different prompts, we can proactively approach reliable LLM services by prompt selection. This case study goes in a similar vein to stolfo2024improving-instruction-following-via-activation-steering; cao2024worst-prompt-performance-of-llms; heo2025do-llms-know-internally-when-they-follow-instructions; heo2025do-instruction-following-uncertainty-post-response, but our focus is to predict the instruction-following prior to generation and across various cousin prompts. We experiment with Qwen3-8B and Qwen2.5-7B-Instruct. For each model, we randomly sample 1,200 of them as the validation set, with 600 followed and 600 unfollowed. We evaluate three prediction methods, with performance presented in Table˜2.
$\bullet$ Verbalized Confidence. We instruct the LLM to articulate its confidence in successfully following one given prompt, on a scale from 0 to 9. Yet, the self-reported confidence yields performance close to random guessing. This outcome can be attributed to the LLMs’ general tendency toward overconfidence, aligning with xiong2024can-llms-express-their-uncertainty.
@@
$\bullet$ Probing Hidden States. Causal LMs may plan their responses before generation, and the relevant signals can be probed from their hidden states (dong2025emergent-planning-pre-generation-probing). Specifically, we train a linear logistic regression probe, which takes the hidden states of the last input token as input. To train the probe, we additionally collect a non-overlapping set of 200 followed and 200 unfollowed prompts. Compared with the previous two methods, probing provides a more effective prediction. The best validation performance is achieved when probing the 17th and the 27th layer on two models, respectively. Although probing shows potential in this direction, the probes are far from being a reliable predictor.
@@ 5.2 Training-Based Method
We adopt supervised fine-tuning (SFT) as our training method and evaluate two dataset configurations, which represent the two extremes of prompt relatedness. (1) We use the general-domain instruction-following dataset, alpaca (alpaca), whose prompts are largely unrelated to IFEval++. (2) We curate an additional set of prompts, also augmented from IFEval but decontaminated from IFEval++ prompts via exact matching. Then, we collect responses via rejection sampling against LLaMA-3.3-70B-Instruct. We fine-tune Qwen-2.5-7B-Instruct for 312 steps on each dataset. Further details on data curation and training are provided in Section˜E.1.
Figure˜5 reports the reliability performance on IFEval++ across training steps. The two datasets lead to markedly different outcomes. Models fine-tuned on alpaca exhibit a slightly declining performance. In contrast, training on the curated cousin prompts steadily improves reliability, with performance climbing above 45% after 200 steps. This reaffirms reliability as a second-order property from the training perspective. It benefits more from targeted fine-tuning on semantically adjacent samples rather than from the sheer scale of training data. This observation resonates with recent explorations on rewriting training data (fujii2025rewriting-pre-training-data-boosts-llm-performance-in-math-and-code; team2025kimi-k2). In Section˜E.2, we additionally explore how different training techniques (SFT vs. DPO (rafailov2024dpo)) affect the nuance-oriented reliability.
@@ 5.3 Test-Time Scaling
We explore two forms of scaling up test-time compute (snell2024scaling-test-time-compute-optimality): sequential and parallel.
$\bullet$ Sequential (Reasoning Effort). Reasoning models inherently capture thinking patterns such as self-correction (guo2025deepseek-r1). Scaling up reasoning effort can be viewed as a sequential analogue to scaling up test-time compute. We include four models that support adaptive reasoning budget control: GPT-5-mini, Gemini-2.5-Flash, Qwen3-14B, and Qwen3-32B. As shown in Figure˜6, increasing reasoning effort usually improves reliability, particularly for GPT-5-mini. Manual inspection of their intrinsic CoT reveals that, through reasoning, models can infer user intent and adapt responses accordingly. This explicit reasoning process enhances robustness against subtle prompt variations. However, inappropriate reasoning patterns may reduce reliability (chen2024not-tencent-llm-overthinking; dong2025towards-lrm-cog-habits), as evidenced by the negative effects observed in Qwen3-32B.
$\bullet$ Parallel (Rejection Sampling). To scale test-time compute in parallel, we adopt sampling rather than greedy decoding, generating nn samples with a temperature of 1. We assume the availability of a response selector. For example, in prompts related to IFEval++, format requirements can be extracted with the assistance of an LLM, while Python code can be used for rapid response evaluation. With the response selector, the reliable@10 metric is scored positively if the 10 cousin prompts within one group all have at least one response that passes among the nn samples. The results are shown in Figure˜7. Rejection sampling with a suitable response selector effectively enhances the reliability of LLMs in instruction-following. The reliable@10 scores rise significantly as more samples are allowed, plateauing around n=12n=12. We further observe that reasoning models, such as Qwen3-8B, benefit more from additional compute via parallel scaling, given their more exploratory nature (chen2025pass-at-k-training-lrm). For instance, Qwen3-4B and Qwen3-8B, with only n=3n=3 samples, already surpass LLaMA-3.3-70B (the strongest open-source model in Table˜1). However, we note that extending this parallel scaling strategy to other domains may be challenging without suitable response selectors.
@@ 8 Ethical Considerations
This research aims to advance the understanding of the reliability of large language models in instruction-following. All experiments were performed on publicly available models and datasets, ensuring compliance with relevant terms of service and data usage policies, e.g., IFEval (zhou2023instruction-ifeval-verification-instruction-following-evaluation-25-types). No proprietary or confidential information was accessed or reverse-engineered during this study. Our analyses do not involve human subjects, sensitive personal data, or the generation of harmful content. Our use of AI assistants is limited to writing polishing and grammar checking. We release the complete codebase and datasets (e.g., IFEval++) to ensure full reproducibility of our results.
@@ A.1 Benchmarking Instruction-Following
Collie-v1 (yao2024collie-constrained-text-generation) 2,080
CELLO (he2024can-large-language-models-understand-real-world-complex-instructions-aaai-cello) 523
NPB (sun2023evaluating-controlled-generation-tasks)
FollowBench (jiang2024followbench-multi-constraints-composition-and-considering-decoding-methods) 820
IFEval (zhou2023instruction-ifeval-verification-instruction-following-evaluation-25-types) 541
CoDI-Eval (chen2024benchmarking-CoDI-Eval-aaai-diversified-instructions) 9,060
InFoBench (qin2024infobench-evaluate-with-constraint-embedded-and-eval-via-binary-questions) 500
FOFO (xia2024fofo-a-benchmark-to-evaluate-llms-format-following-capability-acl-long-papers) 494
AlpacaEval-LI (yuan2024following-length-constraints-in-instructions) 802
MT-Bench-LI (yuan2024following-length-constraints-in-instructions) 240
SIFo (chen2024sifo-sequential-instruction-following) 800
RuleBench (sun2024beyond-inferential-rule-following-rulebench)
ComplexBench (wen2024benchmarking-complexbench-multiple-constraints-composition) 1,150
CFBench (zhang2024cfbench-comprehensive-constraints-following-benchmark-for-llms) 1,000
SysBench (qin2024sysbench-system-message-following-benchmark-with-constraints) 500
MMMT-IF (epstein2024mmmt-if-multimodal-multi-turn-instruction-following) 71
RealInstruct (ferraz2024llm-realinstruct-real-world-instruction-following-emnlp) 302
Multi-IF (he2024multi-if-multi-turn-and-multilingual-instructions-following-meta) 4,501
LIFBench (wu-etal-2025-lifbench-instruction-following-in-long-context) 2,766
VFF (wang2025verifiable-vff-naacl-format-control) 21k
StructFlowBench (li2025structflowbench-structured-flow-benchmark-for-multi-turn-instruction-following) 155
PBIF (zeng2025order-position-bias-in-multi-constraint-instruction-following) 24k
CodeIF (yan2025codeif) 1,200
WildIFEval (lior2025wildifeval) 11,813
XIFBench (li2025xifbench-multilingual-instruction-following) 465
CodeIF-Bench (wang2025codeif-bench-interactive-code-generation)
MulDimIF ye2025multi-dimensional-constraint-framework-fudan 1,200
MathIF (fu2025scalingreasoninglosingcontrol) 420
AGENTIF (qi2025agentifbenchmarkinginstructionfollowing) 707
LIFEBench (zhang2025lifebench-evaluating-length-instruction) 10,800
MaXIFE (liu2025maxifemultilingualcrosslingualinstruction) 18,285
EifBench (zou2025eifbench) 1,000
Ordered CommonGen (song2025ifircomprehensivebenchmarkevaluating) 27,648
IFBench (pyatkin2025generalizing-if-bench) 300
IFScale (jaroslawicz2025instructionsllmsfollowonce)
IFEvalCode (yang2025ifevalcodecontrolledcodegeneration) 1,600
@@
Collie-v1 (yao2024collie-constrained-text-generation). This work proposes a grammar-based framework for constructing controlled text generation, allowing automatic conversion from formal constraint configurations to natural language instructions. The final compiled benchmark includes 2,080 instances comprising 13 constraint structures.
@@
CELLO (he2024can-large-language-models-understand-real-world-complex-instructions-aaai-cello). Accounting for the complexity of in-the-wild user instructions, this work focuses on long and complex instructions. The proposed CELLO benchmark contains 523 instructions that cover four types of response constraints: count limit, answer format, input dependency, and phrase callback.
@@
NPB (sun2023evaluating-controlled-generation-tasks). This work evaluates the controllability of LLMs during generation tasks. It includes numerical planning constraints, requesting responses with varying lengths.
@@
FollowBench (jiang2024followbench-multi-constraints-composition-and-considering-decoding-methods). This work assesses the constraint-following abilities of LLMs, considering five types of constraints: content, situation, style, format, and example. It involves incrementally adding constraints when curating instructions, yielding a total of 820 instructions with different difficulty levels.
@@
IFEval (zhou2023instruction-ifeval-verification-instruction-following-evaluation-25-types). This work pioneers in proposing a broad taxonomy of code-verifiable constraints. The resulting benchmark consists of 541 instructions, covering 25 types of constraints.
@@
CoDI-Eval (chen2024benchmarking-CoDI-Eval-aaai-diversified-instructions). This work introduces expansion and rewriting steps to enhance instruction diversity. The resulting CoDI-Eval benchmark consists of 9,060 instructions and covers a range of constraints, including sentiment, topic, keyword, length, and toxicity avoidance.
@@
InFoBench (qin2024infobench-evaluate-with-constraint-embedded-and-eval-via-binary-questions). This work proposes a novel scoring method: preparing a checklist for each instruction and employing LLM-as-a-Judge to decide the acceptance. Meanwhile, it introduces the InFoBench benchmark, which comprises 500 instructions and involves constraints like content constraint, linguistic constraint, style rules, format specifications, and number limits.
@@
FOFO (xia2024fofo-a-benchmark-to-evaluate-llms-format-following-capability-acl-long-papers). This work specifically studies the format-following abilities of LLMs. It proposes the FoFo benchmark, which contains 494 instructions. Each test case requires the responses to adhere to domain-specific formats, e.g., medical reports and Latex. LLM-as-a-Judge is employed for automated adherence evaluation.
@@
AlpacaEval-LI & MT-Bench-LI (yuan2024following-length-constraints-in-instructions). This work investigates LLMs’ ability to follow length constraints in instructions. To facilitate this evaluation, the authors extend AlpacaEval (dubois2024length-controlled-alpacaeval) and MT-Bench (zheng2024llm-judge-fastchat) by constructing the corresponding length-instructed variants
@@
SIFo (chen2024sifo-sequential-instruction-following). This work focuses on a specific aspect of instruction-following evaluation—ordering of multiple instructions (i.e., constraints in this work). The SIFo benchmark comprises 800 test cases, with LLMs explicitly instructed to address multiple constraints in a prescribed sequential order. This is implemented as a single-turn interaction.
@@
RuleBench (sun2024beyond-inferential-rule-following-rulebench). This work studies inferential rule following, where additional if–then rules are incorporated into the instructions. This setting differs from the affirmatively requested constraints in other works.
@@
ComplexBench (wen2024benchmarking-complexbench-multiple-constraints-composition). This work studies LLMs’ abilities to follow complex instructions involving multiple constraint compositions. The ComplexBench consists of 1,150 instructions, each associated with a constraint composite. The composition types considered include And, Chain, Selection, and Nested.
@@
CFBench (zhang2024cfbench-comprehensive-constraints-following-benchmark-for-llms). This work emphasizes the comprehensiveness of constraint types for evaluating instruction following. It proposes a constraint system with 25 subcategories and meticulously curates 1K Chinese instructions.

### v2

@@ Abstract
Jianshuo Dong1, Yutong Zhang1, Yan Liu2, Zhenyu Zhong2, Tao Wei2, Chao Zhang1, Han Qiu1thanks: Corresponding author. 1Tsinghua University, China. 2Ant Group, China.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable abilities to follow natural language instructions (Ouyang et al., 2022). The instruction-following ability is foundational for enabling faithful user interaction (Zhou et al., 2023; Lior et al., 2025; Zhang et al., 2024b), reliable agent behavior (Zhang et al., 2025a), and reduced harmful outcomes (Ruan et al., 2024). Evaluating this ability is essential for building reliable, trustworthy AI systems. Consequently, numerous benchmarks have been proposed to measure instruction-following performance across diverse task types and constraint categories (see Appendix˜A for a survey of 36 benchmarks).
@@
As illustrated in Figure˜1, the instruction-following ability of LLMs has improved rapidly across generations. Recent LLMs achieve near-ceiling performance on widely adopted benchmarks like IFEval (Zhou et al., 2023), with GPT-5 reaching an accuracy as high as 95.9%. However, growing evidence reveals that model performance can be highly sensitive to prompt wording (Sclar et al., 2024; Mizrahi et al., 2024; Cao et al., 2024), and that overfitting to benchmark test cases is possible (Zhang et al., 2024a). Thus, questions remain about how well benchmark accuracy translates to real-world reliability.
@@
To operationalize this, we propose a new metric, reliable@k, which quantifies whether an LLM can simultaneously solve a set of cousin prompts. We further design an automated pipeline for generating high-quality cousin prompts using three data augmentation strategies: rephrasing, distractor addition, and task/constraint reconfiguration. This is coupled with a code-assisted validity checker. Building on IFEval (Zhou et al., 2023), we leverage the pipeline to construct an extended benchmark, IFEval++ ***We name it to show respect for the renowned IFEval benchmark, as we position IFEval++ as a successor of IFEval, similar to how C++ relates to C. , which consists of 541 test cases, each comprising 10 cousin prompts. This provides a testbed for assessing nuance-oriented reliability of LLMs in a systematic way.
@@ 2.2 Scope of This Work
Data Source & Constraint Types. We choose IFEval (Zhou et al., 2023) as our main testbed, since SOTA LLMs have already achieved strong performance on it (e.g., OpenAI’s GPT-5 reaches an accuracy of 95.9% as measured in Table˜1), making it a meaningful setting for discussing reliability and the challenging reliable@k metric. This benchmark focuses on single-turn tasks with 25 types of code-verifiable format constraints, such as “Answer using at most {num_words} words.” Reliably tackling such simple yet foundational instructions is a necessary condition for addressing more complex, multi-turn, and multi-modal instruction-following settings. See Section˜B.1 for the 25 constraints.
@@ 3.1 Data Augmentation Methods
Type 2. Distractor Addition. The augmenter appends each prompt with a distractor constraint that does not interfere with the satisfaction of the original constraints (see Section˜E.2). The augmenter is instructed to ensure that each distractor constraint remains compatible with the original prompt. In this setting, responses continue to be evaluated against the original constraints, with the added constraint functioning purely as a distractor, in contrast to prior work on multi-constraint composition (Jiang et al., 2024; Wen et al., 2024). This setup examines whether LLMs can integrate additional constraints in a plug-and-play manner.
@@ 3.2 Code-Assisted Validity Check
The checker is designed with an emphasis on high recall: its primary objective is to flag any potentially erroneous cases, even at the cost of over-detection. This strategy allows us to curate a high-quality evaluation set by filtering candidates against the checker. Inspired by Chen et al. (2023), our central design is to embed the implementations of evaluation functions directly into the checker’s prompt, clearly specifying how the evaluation operates (see Section˜E.2). Empirically, this approach enables more deterministic and robust verification. The checker is implemented in a unified way to monitor all three types of augmented test cases.
@@
To further assess robustness, we evaluated it on 3×1,000=3,0003$\times$ 1,000=3,000 additional flawed cases, maintaining a recall of 99.9% (2,997/3,000); the few misses were due to unavoidable hallucinations (Sun et al., 2025). The checker also achieved 90.0% precision on 541 unrevised promptsMany flagged cases were indeed flawed or ambiguous. Necessary corrections are detailed in Section B.3, and the revised dataset is used for subsequent experiments. . Overall, the checker rigorously detects invalid prompts with minimal valid loss. Moreover, since the augmenter is also explicitly guided to avoid introducing flaws, the two-layer mechanism minimizes the risk of flaws arising from the augmentation process.
@@ 4.1 Experimental Setup
Evaluation Methods. Following the practices of DeepSeek-V3 (Liu et al., 2024), Qwen3 (Yang et al., 2025a), and Kimi K2 (Team et al., 2025), we use the prompt strict evaluation mode of IFEval. This requires LLMs to satisfy all requirements specified in the prompts. Based on this success signal, we compute the reliable@k metric.
@@ 4.2 Main Results with IFEval++
$\bullet$ Augmentation Type. Models exhibit varying robustness to the three types of cousin prompts. Most models can effectively handle rephrased cousin prompts, which share the greatest similarity with the original ones. This result is expected, as rephrased prompts differ only in wording (understanding the instructions) while requesting the same underlying response (executing the instructions). In contrast, cousin prompts that include additional distractor constraints or C/T reconfiguration introduce greater complexity. Their corresponding reliable@4 scores are typically lower than those of rephrased prompts. We hypothesize that, although distractor constraints are designed to remain compatible with the original instructions, and C/T reconfiguration introduces only slight variations in the underlying requirements, both inevitably increase the difficulty of response planning and execution (Dong et al., 2025b).
@@ 5.1 Predicting the Selective Following
If we can predict how the LLM reacts to different prompts, we can proactively approach reliable LLM services by prompt selection. This case study goes in a similar vein to Stolfo et al. (2025); Cao et al. (2024); Heo et al. (2025a, b), but our focus is to predict the instruction-following prior to generation and across various cousin prompts. We experiment with Qwen3-8B and Qwen2.5-7B-Instruct. For each model, we randomly sample 1,200 of them as the validation set, with 600 followed and 600 unfollowed. We evaluate three prediction methods, with performance presented in Table˜2.
$\bullet$ Verbalized Confidence. We instruct the LLM to articulate its confidence in successfully following one given prompt, on a scale from 0 to 9. Yet, the self-reported confidence yields performance close to random guessing. This outcome can be attributed to the LLMs’ general tendency toward overconfidence, aligning with Xiong et al. (2024).
@@
$\bullet$ Probing Hidden States. Causal LMs may plan their responses before generation, and the relevant signals can be probed from their hidden states (Dong et al., 2025b). Specifically, we train a linear logistic regression probe, which takes the hidden states of the last input token as input. To train the probe, we additionally collect a non-overlapping set of 200 followed and 200 unfollowed prompts. Compared with the previous two methods, probing provides a more effective prediction. The best validation performance is achieved when probing the 17th and the 27th layer on two models, respectively. Although probing shows potential in this direction, the probes are far from being a reliable predictor.
@@ 5.2 Training-Based Method
We adopt supervised fine-tuning (SFT) as our training method and evaluate two dataset configurations, which represent the two extremes of prompt relatedness. (1) We use the general-domain instruction-following dataset, alpaca (Taori et al., 2023), whose prompts are largely unrelated to IFEval++. (2) We curate an additional set of prompts, also augmented from IFEval but decontaminated from IFEval++ prompts via exact matching. Then, we collect responses via rejection sampling against LLaMA-3.3-70B-Instruct. We fine-tune Qwen-2.5-7B-Instruct for 312 steps on each dataset. Further details on data curation and training are provided in Section˜E.1.
Figure˜5 reports the reliability performance on IFEval++ across training steps. The two datasets lead to markedly different outcomes. Models fine-tuned on alpaca exhibit a slightly declining performance. In contrast, training on the curated cousin prompts steadily improves reliability, with performance climbing above 45% after 200 steps. This reaffirms reliability as a second-order property from the training perspective. It benefits more from targeted fine-tuning on semantically adjacent samples rather than from the sheer scale of training data. This observation resonates with recent explorations on rewriting training data (Fujii et al., 2025; Team et al., 2025). In Section˜E.2, we additionally explore how different training techniques (SFT vs. DPO (Rafailov et al., 2024)) affect the nuance-oriented reliability.
@@ 5.3 Test-Time Scaling
We explore two forms of scaling up test-time compute (Snell et al., 2024): sequential and parallel.
$\bullet$ Sequential (Reasoning Effort). Reasoning models inherently capture thinking patterns such as self-correction (Guo et al., 2025). Scaling up reasoning effort can be viewed as a sequential analogue to scaling up test-time compute. We include four models that support adaptive reasoning budget control: GPT-5-mini, Gemini-2.5-Flash, Qwen3-14B, and Qwen3-32B. As shown in Figure˜6, increasing reasoning effort usually improves reliability, particularly for GPT-5-mini. Manual inspection of their intrinsic CoT reveals that, through reasoning, models can infer user intent and adapt responses accordingly. This explicit reasoning process enhances robustness against subtle prompt variations. However, inappropriate reasoning patterns may reduce reliability (Chen et al., 2024a; Dong et al., 2025a), as evidenced by the negative effects observed in Qwen3-32B.
$\bullet$ Parallel (Rejection Sampling). To scale test-time compute in parallel, we adopt sampling rather than greedy decoding, generating nn samples with a temperature of 1. We assume the availability of a response selector. For example, in prompts related to IFEval++, format requirements can be extracted with the assistance of an LLM, while Python code can be used for rapid response evaluation. With the response selector, the reliable@10 metric is scored positively if the 10 cousin prompts within one group all have at least one response that passes among the nn samples. The results are shown in Figure˜7. Rejection sampling with a suitable response selector effectively enhances the reliability of LLMs in instruction-following. The reliable@10 scores rise significantly as more samples are allowed, plateauing around n=12n=12. We further observe that reasoning models, such as Qwen3-8B, benefit more from additional compute via parallel scaling, given their more exploratory nature (Chen et al., 2025). For instance, Qwen3-4B and Qwen3-8B, with only n=3n=3 samples, already surpass LLaMA-3.3-70B (the strongest open-source model in Table˜1). However, we note that extending this parallel scaling strategy to other domains may be challenging without suitable response selectors.
@@ 6 Conclusion
7 Acknowledgment
This work was supported by Ant Group and the Center for High Performance Computing at Tsinghua University.
@@ 9 Ethical Considerations
This research aims to advance the understanding of the reliability of large language models in instruction-following. All experiments were performed on publicly available models and datasets, ensuring compliance with relevant terms of service and data usage policies, e.g., IFEval (Zhou et al., 2023). No proprietary or confidential information was accessed or reverse-engineered during this study. Our analyses do not involve human subjects, sensitive personal data, or the generation of harmful content. Our use of AI assistants is limited to writing polishing and grammar checking. We release the complete codebase and datasets (e.g., IFEval++) to ensure full reproducibility of our results.
@@ A.1 Benchmarking Instruction-Following
Collie-v1 (Yao et al., 2024) 2,080
CELLO (He et al., 2024a) 523
NPB (Sun et al., 2023)
FollowBench (Jiang et al., 2024) 820
IFEval (Zhou et al., 2023) 541
CoDI-Eval (Chen et al., 2024c) 9,060
InFoBench (Qin et al., 2024) 500
FOFO (Xia et al., 2024) 494
AlpacaEval-LI (Yuan et al., 2024) 802
MT-Bench-LI (Yuan et al., 2024) 240
SIFo (Chen et al., 2024b) 800
RuleBench (Sun et al., 2024)
ComplexBench (Wen et al., 2024) 1,150
CFBench (Zhang et al., 2024b) 1,000
SysBench (Qin et al., 2025) 500
MMMT-IF (Epstein et al., 2024) 71
RealInstruct (Ferraz et al., 2024) 302
Multi-IF (He et al., 2024b) 4,501
LIFBench (Wu et al., 2025) 2,766
VFF (Wang et al., 2025b) 21k
StructFlowBench (Li et al., 2025a) 155
PBIF (Zeng et al., 2025) 24k
CodeIF (Yan et al., 2025) 1,200
WildIFEval (Lior et al., 2025) 11,813
XIFBench (Li et al., 2025b) 465
CodeIF-Bench (Wang et al., 2025a)
MulDimIF Ye et al. (2025) 1,200
MathIF (Fu et al., 2025) 420
AGENTIF (Qi et al., 2025) 707
LIFEBench (Zhang et al., 2025b) 10,800
MaXIFE (Liu et al., 2025) 18,285
EifBench (Zou et al., 2025) 1,000
Ordered CommonGen (Song et al., 2025) 27,648
IFBench (Pyatkin et al., 2025) 300
IFScale (Jaroslawicz et al., 2025)
IFEvalCode (Yang et al., 2025b) 1,600
@@
Collie-v1 (Yao et al., 2024). This work proposes a grammar-based framework for constructing controlled text generation, allowing automatic conversion from formal constraint configurations to natural language instructions. The final compiled benchmark includes 2,080 instances comprising 13 constraint structures.
@@
CELLO (He et al., 2024a). Accounting for the complexity of in-the-wild user instructions, this work focuses on long and complex instructions. The proposed CELLO benchmark contains 523 instructions that cover four types of response constraints: count limit, answer format, input dependency, and phrase callback.
@@
NPB (Sun et al., 2023). This work evaluates the controllability of LLMs during generation tasks. It includes numerical planning constraints, requesting responses with varying lengths.
@@
FollowBench (Jiang et al., 2024). This work assesses the constraint-following abilities of LLMs, considering five types of constraints: content, situation, style, format, and example. It involves incrementally adding constraints when curating instructions, yielding a total of 820 instructions with different difficulty levels.
@@
IFEval (Zhou et al., 2023). This work pioneers in proposing a broad taxonomy of code-verifiable constraints. The resulting benchmark consists of 541 instructions, covering 25 types of constraints.
@@
CoDI-Eval (Chen et al., 2024c). This work introduces expansion and rewriting steps to enhance instruction diversity. The resulting CoDI-Eval benchmark consists of 9,060 instructions and covers a range of constraints, including sentiment, topic, keyword, length, and toxicity avoidance.
@@
InFoBench (Qin et al., 2024). This work proposes a novel scoring method: preparing a checklist for each instruction and employing LLM-as-a-Judge to decide the acceptance. Meanwhile, it introduces the InFoBench benchmark, which comprises 500 instructions and involves constraints like content constraint, linguistic constraint, style rules, format specifications, and number limits.
@@
FOFO (Xia et al., 2024). This work specifically studies the format-following abilities of LLMs. It proposes the FoFo benchmark, which contains 494 instructions. Each test case requires the responses to adhere to domain-specific formats, e.g., medical reports and Latex. LLM-as-a-Judge is employed for automated adherence evaluation.
@@
AlpacaEval-LI & MT-Bench-LI (Yuan et al., 2024). This work investigates LLMs’ ability to follow length constraints in instructions. To facilitate this evaluation, the authors extend AlpacaEval (Dubois et al., 2024) and MT-Bench (Zheng et al., 2023) by constructing the corresponding length-instructed variants
@@
SIFo (Chen et al., 2024b). This work focuses on a specific aspect of instruction-following evaluation—ordering of multiple instructions (i.e., constraints in this work). The SIFo benchmark comprises 800 test cases, with LLMs explicitly instructed to address multiple constraints in a prescribed sequential order. This is implemented as a single-turn interaction.
@@
RuleBench (Sun et al., 2024). This work studies inferential rule following, where additional if–then rules are incorporated into the instructions. This setting differs from the affirmatively requested constraints in other works.
@@
ComplexBench (Wen et al., 2024). This work studies LLMs’ abilities to follow complex instructions involving multiple constraint compositions. The ComplexBench consists of 1,150 instructions, each associated with a constraint composite. The composition types considered include And, Chain, Selection, and Nested.
@@
CFBench (Zhang et al., 2024b). This work emphasizes the comprehensiveness of constraint types for evaluating instruction following. It proposes a constraint system with 25 subcategories and meticulously curates 1K Chinese instructions.